OpenStudy (anonymous):

Add and simplify x/x^2 -9 + 3/x^2-9

OpenStudy (anonymous):

I assume you mean x / (x^2 - 9) + 3/(x^2 - 9) Adding you get (x + 3) / (x^2 - 9) Remember DOTS (Difference Of Two Squares) (x^2 - 9) = (x + 3)(x - 3) And finally (x + 3) / (x^2 - 9) = (x + 3) / (x + 3)(x - 3) = 1 / (x - 3)

OpenStudy (anonymous):

thank you :)